A robot has fallen from a spaceship, as a result he is orbiting a far away planet all by himself. The robot is in an elliptical orbit with rmin = rR and rmax = 25rR /7, rR is the rescue orbit radius. If the robot changes his orbit to a circle of radius rR then he will be rescued. The robot has loose papers with him which he can throw to change the orbit. M = mass of planet, m = mass of robot, m/3 = mass of papers.

1) Before the robot throw the papers, the papers and him are still travelling together. Find vm, speed at perigee. Express answer in terms of G, M and rR.

2) Calculate vR, the speed of the object in a circular orbit, radius rR, around the planet.

3) At perigee, the robot throws the papers in the same direction he is moving to slow himself down. Find vnd, the speed of the papers relative to the robot right after the papers are thrown.

4) Calculate the eccentricity of the orbit of the notes.
Answer #1

(0) We object of mars m around a planet of mass m im know that relocity of elliptical orbit IGM (22 h is V = Where G= gravita(2) We have VR = GM (2 á rmax Hence, VRE GM 14 Pora) 250R Hence, VR= 49GM 400rr Hence, 7 20 GM NR = VR (3) When the robot isfrom the the sum of total energy of the conservation of energy robot and the paper we have + m vo = ² (m + 3) vm² Hence, tvo?(4) We know the velocity at perigee, Vond jam (ite) Now as lore 7 Hences GM YR VB-1) 7.6M 7 GM lite (te) 16 Hence, ite l-e =